HETAV PATEL

Grade 11,

At 0C and 1 atm pressure the densities of air, oxygen and nitrogen are 1.293, 1.429 and 1.251 kg/m3. Calculate the percentage of nitrogen in air, assuming only these two gases to be present.

At 0C and 1 atm pressure the densities of air, oxygen and nitrogen are 1.293, 1.429 and 1.251 kg/m3. Calculate the percentage of nitrogen in air, assuming only these two gases to be present.

Grade:

1 Answers

Chetan Mandayam Nayakar
11 years ago

PV=nRT

P(N)+P(O)=1 atm

P(N)=n(N)RT/V

P(O)=n(O)RT/V

n=PV/RT

nM/V=PM/RT=density

nM=PVM/RT=mass